OpenStudy (anonymous):

t^2(dx/dt)+3tx=t^4lnt+1, x(1)=0 solve initial value problem this is what I got for standard form dx/dt+3tx/t^2=t^4lnt/t^2+1/t2 final standard form dx/dt+3x/t=t^2lnt+1/t^2.......not sure if its right

OpenStudy (anonymous):

my algebra is HORRIBLE :(

zepdrix (zepdrix):

Yah looks good so far :) Were you able to find an integrating factor in order to solve for the general solution?

OpenStudy (anonymous):

no I stopped there because I wasn't sure if I divided out the t^2 right...ill continue it now though

OpenStudy (anonymous):

t^3 is what I got for the I.F.

zepdrix (zepdrix):

ya me too, i think you're on the right track.

OpenStudy (anonymous):

ok getting a little tough from here...not sure what to do

zepdrix (zepdrix):

So you multiply through by your I.F. which gives us,\[\large t^3x'+3t^2x=t^5 \ln t+t\] Don't forget to multiply BOTH sides by t^3 :) Then our left side will simplify down a bit,\[\large (t^3x)'=t^5 \ln t+t\]Did you get that far?

OpenStudy (anonymous):

sure did....glad I did that part right lol

OpenStudy (anonymous):

give me a sec to try and get the next part

OpenStudy (anonymous):

after integrating I got 3xt^2=(t^6/6)lnt+c

zepdrix (zepdrix):

hmm

zepdrix (zepdrix):

the left side should simply be giving us xt^3, dropping the brackets that tell us to take the derivative. I'm a lil worried about your right side.. should have needed to apply Integration by Parts over there, giving you another term or two.

OpenStudy (anonymous):

this is what I tried integrating \[\int\limits_{?}^{?} d/dt (3xt^2)=\int\limits_{?}^{?}t^5lnt+t\]

OpenStudy (anonymous):

oh ok I didn't use integration by parts :(

zepdrix (zepdrix):

We had this, didn't we? \[\large \int\limits (t^3x)'=\int\limits t^5 \ln t+t\;dt\]

OpenStudy (anonymous):

my left side was \[\int\limits t^3x'+3xt^2\] guess I should've left out the 3xt^2 huh

OpenStudy (anonymous):

i tried multiplying the IF which was t^3 with (3x/t) to give me 3xt^2

zepdrix (zepdrix):

When you multiplied the left side by your integrating factor, you are correct, it gave you,\[\large t^3x'+3t^2x\]Which simplifies down to,\[\large (t^3x)'\]Right? :o

zepdrix (zepdrix):

Product Rule in reverse :o

OpenStudy (anonymous):

is that what happens at that point? reverse product rule? no wonder im having trouble

zepdrix (zepdrix):

we multiply through by our integrating factor, and it gives us something of the form,\[\large uv'+u'v\]And yes, then we simplify it down using the product rule in reverse :)\[\large (uv)'\] In our case, see how the 3t^2 is our (t^3)' ?

zepdrix (zepdrix):

maybe that's confusing the way i said that lol

OpenStudy (anonymous):

no that was helpful thank you

zepdrix (zepdrix):

What we really have on the left side is,\[\large (t^3)x'+(t^3)'x\]

zepdrix (zepdrix):

kk c:

OpenStudy (anonymous):

well i attemped it and looked at the answer. IDK this is a tough one what part would i have to do integration by part for?

zepdrix (zepdrix):

The first term on the right requires Parts.\[\large \int\limits t^5 \ln t \;dt\] Let \(\large u=\ln t\) and see if you can get it to work out :D

zepdrix (zepdrix):

\(\large dv=t^5 dt\)

OpenStudy (anonymous):

du i got t^5/6 and v is t^6/6

OpenStudy (anonymous):

well u alone would be 1/t dt

zepdrix (zepdrix):

du alone? ya

OpenStudy (anonymous):

do i have to use u * v - integral vdu?

zepdrix (zepdrix):

ya sounds right.

OpenStudy (anonymous):

ok what i got was lnt (t^6/6)-(t^7/42) (1/t)

zepdrix (zepdrix):

Woops, small boo boo on the second term. You should have cancelled out those t's BEFORE integrating that second part. So your v du gave you,\[\large \frac{1}{6}\int\limits t^6\frac{1}{t}dt\]Something like that, right? So it would simplify to this before performing the actual integration.\[\large \frac{1}{6}\int\limits t^5 dt\]

OpenStudy (anonymous):

yes that looks better

zepdrix (zepdrix):

It'll just affect the fraction coefficient on that second term.

OpenStudy (anonymous):

so its -2t^6/42 now?

OpenStudy (anonymous):

oops make that 7t^6/42

OpenStudy (anonymous):

lnt(7t^6/42)-(2t^6/42)

OpenStudy (anonymous):

though i don't think that its 42 because the answer has 36 in the denominator

zepdrix (zepdrix):

\[\large \frac{1}{6}\int\limits\limits t^5 dt\qquad=\qquad \frac{1}{6}\left(\frac{1}{5+1}t^{5+1}\right)\qquad=\qquad\frac{1}{36}t^6\]

zepdrix (zepdrix):

I'm not sure where you keep getting this 7 from, heh

OpenStudy (anonymous):

guess my v was wrong. i was using t^6/6 for v

OpenStudy (anonymous):

oh but then the 1/t makes it t^5

zepdrix (zepdrix):

ya :3

OpenStudy (anonymous):

:/ geez my bad lol

OpenStudy (anonymous):

lnt(1/36)?????

zepdrix (zepdrix):

MMk let's just start over with this part -_- you're getting confused lol

OpenStudy (anonymous):

sorry :(

zepdrix (zepdrix):

\[\large \int\limits\limits \color{royalblue}{\ln t} \cdot \color{orangered}{t^5 \;dt}\]\[\large \color{royalblue}{u=\ln t}\qquad\qquad\qquad \color{orangered}{dv=t^5dt}\]\[\large \color{royalblue}{du=\frac{1}{t}dt}\qquad\qquad\quad \color{orangered}{v=\frac{1}{6}t^6}\] So our integral gives us,\[\large \color{orangered}{\frac{1}{6}t^6}\color{royalblue}{\ln t}-\int\limits \left(\color{orangered}{\frac16t^6}\right)\color{royalblue}{\frac1t\;dt}\]Which before integrating again, simplifies to,\[\large \color{orangered}{\frac{1}{6}t^6}\color{royalblue}{\ln t}-\frac{1}{6}\int\limits t^5\;dt\]Which gives us,\[\large \frac{1}{6}t^6 \ln t-\frac{1}{36}t^6\]

zepdrix (zepdrix):

For that portion.. blah i guess i didn't need to do all that color.. took me a bit too long lol

zepdrix (zepdrix):

\[\large \int\limits\limits (t^3x)'=\int\limits\limits t^5 \ln t+t\;dt\]We can rewrite this as,\[\large \int\limits\limits (t^3x)'=\int\limits\limits t^5 \ln t \;dt+\int\limits t\;dt\] And so far we've figured out this much,\[\large t^3x=\frac{1}{6}t^6 \ln t-\frac{1}{36}t^6+\int\limits t\;dt\]

OpenStudy (anonymous):

ok i thought that we had to change 1/6 t^6 to work with the -1/36 t^6

OpenStudy (anonymous):

so now we just divide everything by t^3 right?

zepdrix (zepdrix):

well remember, we didn't integrate the +t on the right yet :o

OpenStudy (anonymous):

oh yeah but that is just 1 so it could just be +C

zepdrix (zepdrix):

woops, you accidentally took it's derivative instead of integrating.

OpenStudy (anonymous):

:/ what a noob lol ok so t^2/2+C

zepdrix (zepdrix):

heh

zepdrix (zepdrix):

\[\large t^3x=\frac{1}{6}t^6 \ln t-\frac{1}{36}t^6+\frac{1}{2}t^2+C\] Ok looks good so far

OpenStudy (anonymous):

I really need to start writing the fractions as 1/6 t^6 instead of t^6/6 cause i forget about that 1/6 or fraction to take out of the integration

zepdrix (zepdrix):

yah i like the fractions in front c: whatever works best for you though.

OpenStudy (anonymous):

but from what we have now i just divide the t^3 out of the x and that should be the soln correct?

OpenStudy (anonymous):

well actually i forgot about our initial condition. had to bring my laptop to the charger and left my work behind lol but ill scroll to the top

zepdrix (zepdrix):

XD

zepdrix (zepdrix):

Use the home and end keys if you have those on your keyboard :D nice quick way to get from bottom to top

zepdrix (zepdrix):

cause i can imagine you're probably using a touchpad to scroll all the way up -_- oh boy lol

OpenStudy (anonymous):

is c=-1?

zepdrix (zepdrix):

hmmm

OpenStudy (anonymous):

or 17/36 according to the ans i think lol

zepdrix (zepdrix):

\[\large t^3x=\frac{1}{6}t^6 \ln t-\frac{1}{36}t^6+\frac{1}{2}t^2+C\]Plugging in our initial value,\[\large 1\cdot 0=\frac{1}{6}\cdot 1\cdot \ln 1-\frac{1}{36}\cdot1^6+\frac{1}{2}\cdot1^2+C\]Hmm I'm coming up with something different, your equation look the same?

zepdrix (zepdrix):

ya the answer key is helpful heh.

OpenStudy (anonymous):

well the final soln is \[x=(t^3/6)lnt-t^3/36+1/2t-17/36t^3\]

zepdrix (zepdrix):

Did you plug your 0 and 1 in correctly? You shouldn't get -1 for C D: They should be plugged in as I did up above. If your brain hurting too much at this point I understand XD lol

OpenStudy (anonymous):

yeah i did do it wrong initially but i figured that it was because ln(0) doesn't work

OpenStudy (anonymous):

i just wasn't sure about dividing out the t^3 before using our initial condition

zepdrix (zepdrix):

Yah, before or after :D whichever you prefer.

OpenStudy (anonymous):

yes its correct! thank you so much for bearing with me. this was def the hardest problem he assigned. I can't thank you enough for all your help!!!!!!

zepdrix (zepdrix):

no probs \c:/ try to get a little more practice with integrating factor. that product rule in reverse thing can be tricky to get used to.

OpenStudy (anonymous):

yeah i know i need to review a lot. is there anything i can do to help you on the site besides giving you the best response?

zepdrix (zepdrix):

aw thats sweet XD lol naw i just like helping people :D not looking for anything heh

OpenStudy (anonymous):

just wrote a testimonial and became a fan! thanks again and have a good night

zepdrix (zepdrix):

you too \c: